1. An auditorium has dimensions 10.0 m x 10.0 m x 5.0 m. How many molecules of air fill the auditorium at 30.0°C and a pressure of 101 kPa?
1
Expert's answer
2020-12-31T08:07:27-0500

volume of the auditorium = 10.0m × 10.0m × 5.0m = 500.0 m³

Temperature = 30.0°C = 273K + 30.0°C = 303 K

Pressure = 101kPa = 101 × 10³ Pa = 101000 Nm-3


from the ideal gas law,

"PV = nRT"


"101000 \u00d7 500.0 = n \u00d7 8.314 \u00d7 303"


"n = \dfrac{ 50500000}{2519.142} = 20,046.5" moles



number of molecules = n × 6.02 × 10²³ = 20,046.5 × 6.023 × 10²³ = 1.207 × 10²⁸ molecules.


This means that 1.207 × 10²⁸ molecules of air fill the auditorium at 30.0°C and a pressure of 101 kPa